• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Olson

    Log entry time 08:00:27 on May31,2007

    Entry number 203820

    keyword=Owl shift summary

    Owl Shift Summary:

    M.Olson, V. Sulkosky

    0018: Switching to Q2 = 10 fm-2
    E' = 2.035 GeV (30 MeV increase from previous, did not cycle IAW Makis' instructions)

    ***RHRS settings fixed (P_r = 0.5389, Angle = 65.045 VDC HV = 73%)

    0031: LHRS at new settings (E'=2.0350, Angle=17.6)
    Triggers Downloaded (2035, 1078, He4)
    Requesting 3.0uA
    Run 3914 (He4, Q2=10, L collimator in, 1.25M)
    Run 3915 (Same, 1.25M)

    0045: Dummy target in
    0056: Run 3916 (Dummy, 250k)

    0059: LHRS E' increased by 2% (2057.7GeV) Quads not cycled
    Triggers downloaded
    0100: Run 3917 (He4, Q2=10, LHRS +2%, L collimator in)

    0114: He4 target in, asked for 20uA
    Run 3918 (junk)
    Run 3919 (10uA, He4, Q2=10, LHRS +2%, L collimator in)

    ***Changing RHRS to Q2=39 for later, since we are only doing single arm for Q2=18 and Q2=14.

    0126: RHRS moved to 63.039 degrees in anticipation of Q2=39
    Cycling quads on RHRS
    Set RHRS for Q2=39

    0146: LHRS set for Q2=18 (E'=1.994GeV, Angle=23.75)
    Triggers downloaded
    Run 3920 (23-25uA, He4, Q2=18, L collimator in)

    0155: RHRS set to Q2=39 (P_r=0.5920, Angle=63.039)

    0159: Run 3921 (same as 3920)

    0213: Moving to dummy target
    0220: Dummy target in, requesting 20uA
    Increased to 40uA
    Run 3922 (40uA, Dummy, Q2=18, L collimator in)

    0239: LHRS set for Q2=14 (E'=2.015GeV, Angle=20.851)
    Triggers downloaded
    Current to 25uA
    Run 3923 (25uA, Dummy, Q2=14, L collimator in) *DT=17%*
    0243: Run 3924 (20uA, Dummy, Q2=14, L collimator in) *DT=14%*

    0252: Target to He4

    0259: Run 3925 (5uA, He4, Q2=14, L collimator in) *DT=12%*
    0314: Run 3926 (same as 3925)
    0324: End run 3926, setting LHRS to Q2=39

    0359: Start run 3927 (short run to adjust prescales)
    0405: Start run 3928 (He4, 50uA, Q2=39, collimators OUT)

    0455: Beam down for ~10mins, ending run 3928 (979k)
    0458: Start run 3929

    0557: End run 3929 (He4, 50uA, Q2=39, collimators OUT, 1.25M)
    Start run 3930

    0624: End run 3930 (He4, 50uA, Q2=39, collimiators OUT, 651k)

    0633: Dummy target into beamline, request 40uA
    Start run 3931 (~20 mins)

    0655: End run 3931 (Dummy, 40uA, Q2=39, collimators OUT, 176k)
    Begin configuration change to Q2=29

    0733: Kinematic change to Q2=29 completed (R Dipole was the limiting factor) requesting 40uA for dummy cell run.

    0748: End run 3932 (Dummy, 40uA, Q2=29, collimators OUT, 137k)
    Request target mask to move He4 target into beamline

    0756: He4 target in beamline, requesting 50uA, checking prescales

    0800: Start Run 3933 (He4, 50uA, Q2=29, collimators OUT)